an object start from rest with a constant acceleration 8 m/(s2 ) along a straight line find A- the speed at the end of 5 s B-the average speed for the 5s interval c-the distance traveled
1
Expert's answer
2013-03-13T07:00:25-0400
1) for uniformly accelerated motion
v=v0+at, v0 - initial velocity, a - acceleration, t- time
v=0+8*5=40 (m/s)
2) for uniformly accelerated motion
u=(v0+v1)/2, u - average speed, v0 - initial speed, v1 -final speed
u=(0+40)/2=20 (m/s)
3) for uniformly accelerated motion
s=v0t+at^2/2
s=0*t+8*25/2=100 (m)